DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, November 20, 2002 Padelford C-36 2:30 pm

(Note: Special time!)

New maximum principles for linear elliptic equations

Neil Trudinger (Australian National University)

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.